OpenStudy (anonymous):

Of the 50 U.S. states, 4 have names that start with the letter W. What percentage of U.S. states have names that start with the letter W?

OpenStudy (anonymous):

8%

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

@Exvited if u think its wrong its kool say what u gotta say

OpenStudy (anonymous):

Rosesmarie is correct. 4 of the states start with W. Trying to find the percentage of the U.S states that have names that start with the letter W... (100% is also 1). 4/50 divided by 100% = 2/25 2/25 divided by 100% = 2 divided by 4 = 8%.

OpenStudy (anonymous):

:)